OpenStudy (anonymous):

A bricklayer uses a cylinder to store sand. The cylinder is 5 ft tall and has a radius of 0.5 ft. Each batch of mortar contains 0.48 ft3 of sand. How many batches of mortar can the bricklayer make with 1 cylinder of sand? Use 3.14 to approximate pi, andround your final answer to the nearest whole number. ______

OpenStudy (anonymous):

@shamim

OpenStudy (anonymous):

can u help? @shamim

OpenStudy (anonymous):

@Michele_Laino can u help?

OpenStudy (michele_laino):

ok! I try!

OpenStudy (anonymous):

ok

OpenStudy (michele_laino):

we have to compute the volume of the cylinder first

OpenStudy (anonymous):

ok

OpenStudy (michele_laino):

hint: the volume V of a cylinder is given by the subsequent formula: \[\Large V = \pi \times {R^2} \times H\]

OpenStudy (michele_laino):

where R is the radius, and H is the height of our cylinder

OpenStudy (anonymous):

idk

OpenStudy (michele_laino):

hint: \[\Large V = \pi \times {R^2} \times H = 3.14 \times {\left( {0.5} \right)^2} \times 5 = ...?\]

OpenStudy (michele_laino):

what is V?

OpenStudy (anonymous):

omg i no it but i forgot :c

OpenStudy (anonymous):

srry

OpenStudy (anonymous):

idk

OpenStudy (michele_laino):

we have: \[\Large \begin{gathered} V = \pi \times {R^2} \times H = 3.14 \times {\left( {0.5} \right)^2} \times 5 = \hfill \\ = 3.14 \times 0.25 \times 5 = 3.14 \times 1.25 = 3.925\;f{t^3} \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

so its 3.925?

OpenStudy (anonymous):

is that corrected @Michele_Laino

OpenStudy (anonymous):

wait is it 5?

OpenStudy (michele_laino):

that's the volume of the cylinder. Now we have to divide 3.925 by 0.48: \[\Large {\text{number of butch of mortar}} = \frac{{3.925}}{{0.48}} = ...?\]

OpenStudy (michele_laino):

butches*

OpenStudy (michele_laino):

what do you get?

OpenStudy (anonymous):

ummm?

OpenStudy (michele_laino):

more explanation: in order to make one butch of mortar, we need of 0.48 ft^3 of sand, now our cylinder contains 3.925 ft^3 of sand, so, in order to know how many butches of mortar, we can make with the sand contained in our cylinder, we have to divide 3.925 by 0.48

OpenStudy (anonymous):

im geting really long numbers

OpenStudy (michele_laino):

we have to round off the result to the nearest unit

OpenStudy (anonymous):

oh ok

OpenStudy (michele_laino):

so, what is: \[\Large \frac{{3.925}}{{0.48}} = ...?\] you can use Windows calculator, for example

OpenStudy (anonymous):

8.17708333 ?

OpenStudy (michele_laino):

ok! Now you have to round off that result to the nearest unit, what do you get?

OpenStudy (anonymous):

idk my sis is confusing me shes telling me something else

OpenStudy (anonymous):

can u tell me?

OpenStudy (michele_laino):

the answer is 8

OpenStudy (anonymous):

oooh thats what i thought but my sis said no its not lol

OpenStudy (anonymous):

ty

OpenStudy (michele_laino):

thank you!